Joe invested a portion of $6000 at 4.5% interest and the balance at 6% interest. How much did he invest at each rate if his total income from both investments is $279?

Answer by checkley77(12569)

.06x+.045(6,000-x)=279
.06x+270-.045x=279
.015x=279-270
x=9/.015
x=600
6,000-600=5,400
.06*600+.045*5,400=279
36+243=279
279=279